A box contains two black balls and three gold balls. Two
balls are randomly drawn in succession from the box.
a. If there is no replacement, what is the probability
that both balls are black?
b. If there is replacement before the second draw, what
is the probability that both balls are black?

Answer by jim_thompson5910(35256)

There are 2 black + 3 gold = 5 balls total

a)

P(Both Black) = P(Black)*P(Black) = (2/5)*(1/4) = 2/20 = 1/10

b)

P(Both Black) = P(Black)*P(Black) = (2/5)*(2/5) = 4/25